# Large Language Models as Span Annotators

### v1

@@ Abstract
For high-quality texts, single-score metrics seldom provide actionable feedback. In contrast, span annotation—pointing out issues in the text by annotating their spans—can guide improvements and provide insights. Until recently, span annotation was limited to human annotators or fine-tuned encoder models. In this study, we automate span annotation with large language models (LLMs). We compare expert or skilled crowdworker annotators with open and proprietary LLMs on three tasks: data-to-text generation evaluation, machine translation evaluation, and propaganda detection in human-written texts. In our experiments, we show that LLMs as span annotators are straightforward to implement and notably more cost-efficient than human annotators. The LLMs achieve moderate agreement with skilled human annotators, in some scenarios comparable to the average agreement among the annotators themselves. Qualitative analysis shows that reasoning models outperform their instruction-tuned counterparts and provide more valid explanations for annotations. We release the dataset of more than 40k model and human annotations for further research. [^1] Project website: https://llm-span-annotators.github.io
@@ 1 Introduction
Fine-grained aspects of texts, such as faithfulness or coherence, depend on local lexical choices. We need a detailed analysis in order to reflect them in the evaluation of natural language generation (NLG) systems or quality judgments on human-written texts. However, most automatic evaluation metrics are based on direct assessment, producing holistic scores for each evaluated aspect Gkatzia & Mahamood (2015); Sai et al. (2023); Schmidtová et al. (2024). Although numerical values make it easy to rank systems, these metrics are too simplistic and susceptible to biases or miscalibration issues of the underlying models Gehrmann et al. (2023); Gao et al. (2024); Liu et al. (2024); Wang et al. (2024).
@@
The subject of our study, span annotation, offers an alternative approach. Figure 1 shows the difference between direct assessment and span annotation. Instead of assigning a single score for each evaluated aspect, the goal of span annotation is to localize text spans and classify them into predefined categories. In contrast to numerical rating, annotations are aligned to the evaluated text, which makes them more explainable and actionable. The annotations can also be examined post hoc, enabling a more modular evaluation process.
Despite its advantages, span annotation has not yet been widely applied in automatic NLG evaluation. The method has traditionally relied on human annotators, making it costly and difficult to scale Da San Martino et al. (2019); Thomson & Reiter (2020); Popovic (2020); Kocmi et al. (2024b).
@@
The situation has changed with LLM-as-a-judge paradigm, in which LLMs follow instructions for the evaluation process similarly to human annotators Zheng et al. (2023); Gu et al. (2024). However, the current LLM-as-a-judge applications still mostly focus on document-level classification or numerical ratings on the Likert scale, which are too coarse and unreliable for evaluating fine-grained aspects of texts.
Here, we investigate a specific LLM-as-a-judge application: using LLMs as span annotators. Instructing LLMs to annotate text spans with predefined categories is a way to produce span annotations automatically and robustly, lowering the barrier to using the method for NLG evaluation.
We focus on the following research questions:
  • RQ1: How to instantiate an LLM-based span annotator? We do a small-scale study of prompting techniques to arrive at a setup that is robust across diverse tasks and scenarios.
  • RQ2: How do LLMs compare to skilled human annotators? We compute inter-annotator agreement (IAA) between annotations produced by LLMs and skilled human annotators on both novel data and existing datasets.
  • RQ3: How valid are LLM annotations and their explanations? We manually analyze a subset of LLM and human annotations to assess the validity of annotations and their respective explanations.
For our experiments, we select three diverse tasks (cf. Section 3.1): evaluation of data-to-text generation Thomson & Reiter (2020), evaluation of machine translation Kocmi et al. (2024b), and propaganda detection Da San Martino et al. (2019). We base our span annotators on both open and proprietary state-of-the-art LLMs, including recent reasoning models. We collect more than 40k model and human annotations that we release for further investigation.
In the experimental part, we first analyze various prompting styles, demonstrating that zero-shot prompting with detailed guidelines is a robust way to set up LLMs as span annotators. Using automatic metrics, we show that LLMs are close to IAA of skilled human annotators, leveling or surpassing the performance of qualified crowdworkers in some scenarios. Qualitatively, we find that LLM annotations are completely correct in 49.5% of cases (56.5% in the case of reasoning models), which is similar to qualified crowdworkers.
Figure 1: Comparison between direct assesment of text (=conventional evaluation methods) and span annotation with pre-defined categories (=this paper). Span annotation provides more signal for diagnosing model outputs.
@@ 2 Related Work
Automatic NLG metrics tradionally assess text quality by measuring similarity to human-written reference texts Sai et al. (2023); Schmidtová et al. (2024). As such, they are unable to quantify more fine-grained aspects (Gehrmann et al., 2023; Freitag et al., 2021) and do not correlate well with human judgements Novikova et al. (2017); Reiter (2018). With the emerging LLM-as-a-judge paradigm (Gu et al., 2024), LLMs have been applied as evaluators across diverse tasks, either with simple numeric scoring (Bavaresco et al., 2024; Liu et al., 2023; Sottana et al., 2023; Leiter et al., 2023; Chiang & Lee, 2023) or with detailed feedback (Li et al., 2024; Kim et al., 2024a; b; Xu et al., 2023). However, the scores produced by LLMs still miss fine-grained text aspects and are influenced by LLM biases Stureborg et al. (2024); Koo et al. (2024); Wang et al. (2024).
@@
In machine translation (MT), span annotation is a long-standing component of protocols such as MQM or ESA (Lommel et al., 2014; Mariana, 2014; Popovic, 2020; Kocmi et al., 2024b), where annotators mark erroneous spans in translations. In data-to-text (D2T) generation, span annotation was applied by Thomson & Reiter (2020), who introduce a span-annotation-based evaluation protocol for annotation of generated basketball reports. Span annotation is also used to judge intrinsic text qualities, such as coherence or use of rhetorical devices, in tasks such as propaganda detection Da San Martino et al. (2019), and text summarization Subbiah et al. (2024). Unlike our work, these works conduct only human annotation.
@@
For all the aforementioned tasks, first attempts at automating span annotation were based on fine-tuned pre-trained encoder models. That includes MT (Guerreiro et al., 2024), D2T generation Kasner et al. (2021), text summarization Goyal et al. (2022), and propaganda detection Martino et al. (2020); Goffredo et al. (2023); Piskorski et al. (2023). Automating span annotation with LLMs is more flexible and benefits from increasing LLM capabilities. While existing works (Kocmi & Federmann, 2023; Fernandes et al., 2023; Hasanain et al., 2024; Kasner & Dušek, 2024; Chang et al., 2024; Kartáč et al., 2025; Ramponi et al., 2025; Zouhar et al., 2024) applied LLMs as a tool for evaluating their systems, our study systematically compares LLMs to human annotators across tasks and domains.
See more discussion on related work in Appendix H.
@@ 3.1 Span Annotation: Task Definition
The goal of span annotation is to annotate a text sequence $Y=\langle y_{1},\ldots,y_{n}\rangle$, i.e., to produce a set of annotations $A=\{a_{1},\ldots,a_{m}\}$, where each annotation $a_{i}$ is a tuple $\langle s_{i},e_{i},c_{i}\rangle$:
  • $s_{i},e_{i}\in\{1,\ldots,n\},s_{i}<e_{i}$ are the start and end indices of the annotated span, [^2] Our definition allows overlapping annotations: $s_{i}\leq s_{j}\leq e_{i};i\neq j$.
  • $c_{i}\in C$ is the assigned annotation category from the pre-defined set of task-specific categories $C=\{c_{1},\ldots,c_{k}\}$.
The process is further specified in the guidelines $\mathcal{G}$: a set of fine-grained instructions for the annotation process regarding how to handle ambiguities or how to decide which parts of the text to include in the span. The annotator also considers the source $X$ (e.g., the translation source or the input data). Note that $X$ can be empty if we are annotating only intrinsic text aspects such as coherence or style.
3.2 Automating Span Annotations
Following our definition from Section 3.1, our goal is to collect the set of annotations $A$ for the given input $\langle Y,C,\mathcal{G},X\rangle$. In our setup, we rely on the LLM-as-judge paradigm: prompting an LLM to follow annotation guidelines:
@@ 3.2 Automating Span Annotations
To obtain the annotations, we closely follow the setup of Kasner & Dušek (2024). Specifically, we use structured decoding, asking for the list of annotations in JSON format. The output contains the fields reason (a short sentence justifying the annotation), text (the literal content of the matched span), and type (the integer index of the error category as given in the prompt). This approach enables us to robustly parse the model output and collect auxiliary explanations for annotations. [^3] See Appendix B for a detailed discussion of the selected approach and alternatives.
@@ 3.3 Evaluating Span Annotations
To determine and compare the quality of the human and the automatic annotation process, we need a notion of similarity between two sets of annotations $\{A_{1},A_{2},\ldots,A_{|Y|}\}$ and $\{\hat{A}_{1},\hat{A}_{2},\ldots,\hat{A}_{|Y|}\}$ over a set of texts $Y=\{Y_{1},Y_{2},\ldots,Y_{|Y|}\}$. Based on these similarities, we can then compute reliability metrics, such as IAA or comparison to gold annotations. We consider three similarity metrics.
@@
The simplest correlation computation simply compares how many spans were annotated for each example.
$\mathrm{CountCorrelation}(A_{*},\hat{A}_{*})=\mathrm{Pearson}(\langle|A_{y}|,|\hat{A}_{y}|\rangle_{y\in Y})$ (2)
@@
To also quantify the degree of alignment between individual annotations, we compute precision, recall, and $F_{1}$ as defined in Da San Martino et al. (2019). These measures are on matching annotations, adjusted to give partial credit to imperfect matches (which is necessary given the varying lengths of annotations): $\text{Precision}(A_{*},\hat{A}_{*})=\frac{1}{|A|}\sum_{a\in A}\max_{\hat{a}\in\hat{A}}\frac{|a\cap\hat{a}|}{|a|},$ (3) $\text{Recall}(A_{*},\hat{A}_{*})=\frac{1}{|\hat{A}|}\sum_{\hat{a}\in\hat{A}}\max_{a\in A}\frac{|a\cap\hat{a}|}{|\hat{a}|},$ (4) where $a\cap\hat{a}$ is the character overlap between two annotation spans and $|a|=e-s+1$ is the length of the annotation span in characters (see Section 3.1). Subsequently, we compute $F_{1}$-score as the harmonic mean of precision and recall.
For each of the metrics, we consider soft and hard variants. While the soft variant disregards span categories, the hard one only considers overlaps where the span category is matching. We consider the hard variant to be the default. Additionally, we report the difference $F_{1}\Delta=F_{1}(\text{soft})-F_{1}(\text{hard})$.
@@
The $F_{1}$ score is sensitive to varying span granularities and does not consider near matches with no overlap. It also does not account for agreement by chance. To this end, we follow Da San Martino et al. (2019) and Hasanain et al. (2024) in using the $\gamma$ score (Mathet et al., 2015) as another annotation similarity metric. This metric operates on whole sets of annotations for a single input $y$: $A_{*,y}=\{A,\hat{A}_{y},\ldots,A_{k,y}\}$. The metric builds the best possible alignment between the annotations $A_{i,y}$ and $A_{j,y}$ and computes their dissimilarity $\mathrm{Dissimilarity}(A_{i,y},A_{j,y})$. This is then compared to the expected dissimilarity by chance, which is created by sampling across all annotations: $\mathrm{DissimilarityChance}(A_{*,y})$. The final score is based on the average dissimilarity between annotations and the dissimilarity by chance:
$\gamma(A_{*,y})=1-\frac{\frac{1}{k^{2}}\sum_{i,j=1}^{k}\mathrm{Dissimilarity}(A_{i,y},A_{j,y})}{\mathrm{DissimilarityChance}(A_{*,y})}$ (5)
The score ranges from negative infinity to 1, where 1 is achieved when the annotations have perfect alignment. The $\gamma$ score extends Krippendorff’s $\alpha$ (Krippendorff, 1980), another popular metric, by computing the category-aware span alignments. We use the implementation of Titeux & Riad (2021).
Examples with empty annotation sets.
For an output $Y$, either one or both of the annotation sets $A_{y},\hat{A}_{y}$ can be empty. Note that this case is natural and may even be desirable: e.g., if the goal is to annotate errors in an output that is entirely correct. However, neither of the introduced scores accounts for these cases accordingly. [^4] The $\gamma$ score is undefined for less than two non-empty annotation sets: these examples thus need to be skipped during the $\gamma$ computation. The F1 score does not account for examples with no annotations. To compensate for that, we introduce a score $S_{\emptyset}$ that is computed for examples with less than two non-empty sets of annotations as follows:
$S_{\emptyset}=1/(1+|A|)$ (6)
where $|A|$ is the number of annotations from the remaining annotator. The score is equal to 1 for the cases where no annotator produced any annotations, i.e., a perfect match, and decreases proportionally to the number of annotations from the remaining annotator.

### v2

@@ Abstract
Span annotation is the task of localizing and classifying text spans according to custom guidelines. Annotated spans can be used to analyze and evaluate high-quality texts for which single-score metrics fail to provide actionable feedback. Until recently, span annotation was limited to human annotators or fine-tuned models. In this study, we show that large language models (LLMs) can serve as flexible and cost-effective span annotation backbones. To demonstrate their utility, we compare LLMs to skilled human annotators on three diverse span annotation tasks: evaluating data-to-text generation, identifying translation errors, and detecting propaganda techniques. We demonstrate that LLMs achieve inter-annotator agreement (IAA) comparable to human annotators at a fraction of a cost per output annotation. We also manually analyze model outputs, finding that LLMs make errors at a similar rate to human annotators. We release the dataset of more than 40k model and human annotations for further research. [^1] Project website: https://llm-span-annotators.github.io
@@ 1 Introduction
Fine-grained aspects of texts, such as faithfulness or coherence, depend on local lexical choices. In order to reflect these aspects in quality judgments of texts, we need analysis techniques that provide the appropriate amount of detail. However, most automatic evaluation metrics for natural language generation (NLG) are based on direct assessment, producing holistic scores for each evaluated aspect Gkatzia & Mahamood (2015); Sai et al. (2023); Schmidtová et al. (2024). Although numerical values make it easy to rank systems, these metrics are too simplistic and susceptible to biases or miscalibration of the underlying models Gehrmann et al. (2023); Gao et al. (2024); Liu et al. (2024); Wang et al. (2024).
@@
The subject of our study, span annotation (Figure 1), offers an alternative approach. Instead of assigning a single score for each evaluated aspect, the goal of span annotation is to localize text spans and classify them according to task-specific guidelines. In contrast to numerical rating, annotations are aligned to the evaluated text, which makes them more explainable and actionable. The annotations can also be examined post hoc, which brings more accountability into the evaluation process.
@@
Despite its advantages, span annotation has not yet been widely applied in automatic NLG evaluation. The method traditionally required human annotators, making it costly and difficult to scale Da San Martino et al. (2019); Thomson & Reiter (2020); Popovic (2020); Kocmi et al. (2024c). The situation may change with the upcoming LLM-as-a-judge paradigm, in which LLMs follow instructions for the evaluation process similarly to human annotators. However, current LLM-as-a-judge applications still focus mainly on document-level classification or numerical ratings on the Likert scale Zheng et al. (2023); Gu et al. (2024).
In this work, we investigate whether LLMs can be applied as a substitute for human annotators on span annotation tasks. First, we focus on how to produce span annotations with LLMs. We discuss input tagging methods and experiment with prompting techniques, looking for a setup that is robust and widely applicable. Next, we compare LLMs with skilled human annotators in terms of IAA. Where possible, we contrast the IAA between LLMs and human annotators with the upper bound of the IAA between the annotators themselves. Finally, we manually analyze a subset of LLM and human annotations to assess the validity of the annotations and their respective explanations.
For our experiments, we select three diverse span annotation tasks (cf. Section 3.1): evaluating data-to-text generation Thomson & Reiter (2020), identifying errors in machine translation Kocmi et al. (2024a), and detecting propaganda techniques in human-written texts Da San Martino et al. (2019). Note that unlike established span annotation tasks such as named entity recognition Tjong Kim Sang & De Meulder (2003); Wang et al. (2025a), there are no general purpose fine-tuned models for these tasks since annotation guidelines are specified ad hoc, making their automation difficult.
We base our span annotators on open and proprietary state-of-the-art LLMs, including recent reasoning LLMs that leverage test-time compute for better performance on challenging tasks Muennighoff et al. (2025); Marjanović et al. (2025).
Our contributions are as follows:
  1. 1.
    We demonstrate that zero-shot prompting in combination with structured outputs can be a robust way to set up LLMs as span annotators, given that the model is provided detailed annotation guidelines (Sections 3.2 and 5.1)
  2. 2.
    Using a combination of specialized IAA metrics and manual analysis, we show that performance of LLMs is close to skilled human annotators, in some scenarios leveling or surpassing the performance of qualified crowdworkers (Sections 5.2 and 5.3).
  3. 3.
    We release a dataset of more than 40k human and model annotations for all our tasks and setups, including reasoning traces from reasoning LLMs and annotations collected from skilled crowdworkers.
Figure 1: Examples of span annotation tasks that we automate with LLMs. We unify the setup for evaluation tasks (D2T-Eval, MT-Eval) and text analysis tasks (Propaganda).
@@ 2 Related Work
Automatic NLG metrics traditionally assess text quality by measuring similarity to human-written reference texts Sai et al. (2023); Schmidtová et al. (2024). As such, they are unable to quantify more fine-grained aspects (Gehrmann et al., 2023; Freitag et al., 2021) and do not correlate well with human judgments Novikova et al. (2017); Reiter (2018). With the emerging LLM-as-a-judge paradigm (Gu et al., 2024), LLMs have been applied as evaluators across diverse tasks, either with simple numeric scoring (Bavaresco et al., 2024; Liu et al., 2023; Sottana et al., 2023; Leiter et al., 2023; Chiang & Lee, 2023) or with detailed feedback (Li et al., 2024; Kim et al., 2024a; b; Xu et al., 2023). However, the scores produced by LLMs still miss fine-grained text aspects and are influenced by LLM biases Stureborg et al. (2024); Koo et al. (2024); Wang et al. (2024).
@@
In machine translation (MT), span annotation is a long-standing component of protocols such as MQM or ESA (Lommel et al., 2014; Mariana, 2014; Popovic, 2020; Kocmi et al., 2024c), where annotators mark erroneous spans in translations. In data-to-text (D2T) generation, span annotation is applied by Thomson & Reiter (2020), who introduce a span-based evaluation protocol for annotation of generated basketball reports. Span annotation is also used to judge intrinsic text qualities, such as coherence or use of rhetorical devices, in tasks such as propaganda detection Da San Martino et al. (2019) and text summarization Subbiah et al. (2024). Note that unlike our work, these works perform span annotation only using human annotators.
@@
For all the aforementioned tasks, first attempts at automating span annotation were based on fine-tuned pre-trained encoder models. That includes MT (Guerreiro et al., 2024), D2T generation Kasner et al. (2021), text summarization Goyal et al. (2022), and propaganda detection Martino et al. (2020); Goffredo et al. (2023); Piskorski et al. (2023). Automating span annotation with LLMs is more flexible and benefits from increasing LLM capabilities. We build on recent work that applies LLMs as a task-specific evaluation tool (Kocmi & Federmann, 2023; Fernandes et al., 2023; Hasanain et al., 2024; Kasner & Dušek, 2024; Chang et al., 2024; Zouhar et al., 2025; Kartác et al., 2025; Ramponi et al., 2025). Our work is the first that unifies diverse span annotation tasks within a single framework and systematically compares the performance of LLMs to human annotators.
See more discussion on related work in Appendix F.
@@ 3.1 Span Annotation: Task Definition
The goal of span annotation is to annotate a text sequence $Y=\langle y_{1},\ldots,y_{n}\rangle$ given:
  • the set of categories $C=\{c_{1},\ldots,c_{k}\}$,
  • the annotation guidelines $\mathcal{G}$ (including category definitions and other instructions),
  • the source $X$ (such as the translation source; empty if we are annotating only intrinsic text aspects).
The goal is to produce a set of annotations $A=\{a_{1},\ldots,a_{m}\}$, where each annotation $a_{i}$ is a tuple $\langle s_{i},e_{i},c_{i},r_{i}\rangle$:
  • $s_{i},e_{i}\in\{1,\ldots,n\},s_{i}<e_{i}$ are the start and end indices of the annotated span,
  • $c_{i}\in C$ is the assigned annotation category,
  • $r_{i}$ is an explanation for the annotation (optional). [^2] We collect the explanations from LLMs, but not from human annotators due to time considerations.
3.2 Automating Span Annotation with LLMs
To automate span annotation, we collect the annotations $A$ for the given input $\langle Y,C,\mathcal{G},X\rangle$ from an LLM:
@@ 3.2 Automating Span Annotation with LLMs
How to Identify the Spans?
An annotation needs to contain the start and end indices of the span. For encoder models, the task would be trivially solved by classifying tokens on top of the encoder’s hidden states. However, to the best of our knowledge, there is no established way of tagging the input text with decoder-only models. Of the methods from previous work, we discarded surrounding spans with special XML-like tokens Wang et al. (2025b); Treviso et al. (2024). This method requires the LLM to generate full copy of the input text and does not offer a fallback if the model fails to do so. We also rule out listing start and end indices of spans Hasanain et al. (2024); Ramponi et al. (2025), as Hasanain et al. note that the model frequently generates indices that do not match the corresponding spans in a paragraph. Instead, we adopted the method of listing textual content of the spans Kocmi & Federmann (2023); Zouhar et al. (2025); Kasner & Dušek (2024). Here, the start and end indices of the span are found heuristically by string matching of the span’s textual content. With this approach, we do not need to modify the input and failing to match a single annotation does not influence other matches. [^3] We do note that this approach can misalign the model output by assigning the annotation to the first match, but our preliminary experiments showed that the collisions are rare.
Parsing Model Output
To obtain the annotations, we request the list of annotations in JSON format, following the setup of Kasner & Dušek (2024). This setup allows us to use constrained decoding with a defined JSON scheme, ensuring that the output is syntactically valid. We require each annotation to contain the fields reason (the explanation $r_{i}$), text (the textual content of the span), and type (the integer index of the error category $c_{i}$). [^4] Given the considerations of Castillo (2024), we ensure that the reason field has precedence.
@@ 3.3 Evaluating Span Annotations
To compare annotations automatically, we need a notion of similarity between two sets of annotations $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{|Y|}\}$ and $\hat{\mathcal{A}}=\{\hat{A}_{1},\hat{A}_{2},\ldots,\hat{A}_{|Y|}\}$ over a set of texts $\mathcal{Y}=\{Y_{1},Y_{2},\ldots,Y_{|Y|}\}$. We consider three similarity metrics:
@@
This metric simply compares how many spans were annotated for each example:
$\mathrm{CountCorrelation}(\mathcal{A},\hat{\mathcal{A}})=\mathrm{Pearson}(\langle|A_{Y}|,|\hat{A}_{Y}|\rangle_{Y\in\mathcal{Y}})$ (2)
This metric serves as a sanity check: a low value would suggest that an annotator either skips examples or over-annotates, indicating unclear annotation guidelines.
@@
To quantify the degree of alignment between individual annotations, we compute precision, recall, and $F_{1}$ as defined in Da San Martino et al. (2019). These measures are on matching annotations, adjusted to give partial credit to imperfect matches: $\text{Precision}(A_{Y},\hat{A}_{Y})=\frac{1}{|A_{Y}|}\sum_{a\in A_{Y}}\frac{|a\cap\hat{a}|}{|a|},$ (3) $\text{Recall}(A_{Y},\hat{A}_{Y})=\frac{1}{|\hat{A}_{Y}|}\sum_{\hat{a}\in\hat{A}_{Y}}\frac{|a\cap\hat{a}|}{|\hat{a}|},$ (4) where $a\cap\hat{a}$ is the character overlap between two annotation spans and $|a|=e-s+1$ is the length of the annotation span in characters (see Section 3.1). Subsequently, we compute the $F_{1}$-score as the harmonic mean of precision and recall.
For each of the metrics, we consider soft and hard variants. The hard variant only considers overlaps where the span category matches, while the soft variant disregards the categories. We consider the hard variant to be the default. In addition, we report the difference $F_{1}\Delta=F_{1}(\text{soft})-F_{1}(\text{hard})$.
@@
The $F_{1}$ score is sensitive to varying span granularities and does not consider near-matches with no overlap or agreement by chance. To this end, we follow Da San Martino et al. (2019) and Hasanain et al. (2024) in using the $\gamma$ score (Mathet et al., 2015) as a complementary metric. In short, the metric builds the best possible alignment between the sets of annotations $A_{Y}$ and $\hat{A}_{Y}$ and computes the “disorder” of this alignment based on the positional and categorical dissimilarities of aligned annotations. The score ranges from negative infinity to 1, where 1 is achieved when the annotations are perfectly aligned. The $\gamma$ score extends Krippendorff’s $\alpha$ (Krippendorff, 1980), another popular metric, by computing the category-aware span alignments. We use the implementation of Titeux & Riad (2021).
$S_{\emptyset}$ score.
For an output $y$, one or both annotation sets $A_{Y},\hat{A}_{Y}$ can be empty. Having these cases is desirable: e.g. if the goal is to annotate errors in an output that is entirely correct. However, neither of the introduced scores accounts for these cases. [^5] The $\gamma$ score is undefined for less than two non-empty annotation sets: these examples thus need to be skipped during the $\gamma$ computation. The F1 score does not account for examples with no annotations. To compensate for that, we introduce a score $S_{\emptyset}$ that is computed for examples with less than two nonempty sets of annotations as $S_{\emptyset}=1/(1+|A|)$, where:
$|A|=\begin{cases}|A_{Y}|&\text{if }|\hat{A}_{Y}|=0,\\ |\hat{A}_{Y}|&\text{otherwise}.\end{cases}$ (5)
The score is equal to 1 for the cases where no annotator produced any annotation (i.e., a perfect match) and decreases proportionally to the number of annotations from the annotator that provided a non-zero number of annotations.